### v1

# Are generative models fair? A study of racial bias in dermatological image generation

Abstract
Racial bias in medicine, particularly in dermatology, presents significant ethical and clinical challenges. It often results from the underrepresentation of darker skin tones in training datasets for machine learning models. While efforts to address bias in dermatology have focused on improving dataset diversity and mitigating disparities in discriminative models, the impact of racial bias on generative models remains underexplored. Generative models, such as Variational Autoencoders (VAEs), are increasingly used in healthcare applications, yet their fairness across diverse skin tones is currently not well understood. In this study, we evaluate the fairness of generative models in clinical dermatology with respect to racial bias. For this purpose, we first train a VAE with a perceptual loss to generate and reconstruct high-quality skin images across different skin tones. We utilize the Fitzpatrick17k dataset to examine how racial bias influences the representation and performance of these models. Our findings indicate that the VAE is influenced by the diversity of skin tones in the training dataset, with better performance observed for lighter skin tones. Additionally, the uncertainty estimates produced by the VAE are ineffective in assessing the model’s fairness. These results highlight the need for improved uncertainty quantification mechanisms to detect and address racial bias in generative models for trustworthy healthcare technologies.
Keywords:
AI in medicine Fairness Dermatology Racial bias
1 Introduction
Racial bias in medicine has been widely documented [14, 2]. This bias is often inherited (or even amplified) by deep learning methods, as they lack awareness of sensitive attribute labels [15, 3]. Such limitations can hinder the performance of these algorithms in specific subpopulations, particularly among underrepresented groups. Deploying AI-driven tools in healthcare that may exhibit detrimental performance or consequences for these subgroups is not only unfair but also poses serious risks. For example, an established system that achieves high accuracy may experience an unexpected and dramatic drop in performance when deployed and tested on a different subgroup of the population. Therefore, it is crucial to detect and mitigate these biases to ensure ethical and trustworthy AI in medicine [10].
In dermatology, this racial bias is undeniable due to the significant underrepresentation of various skin tones in clinical data. Despite the promising machine learning results in accurately classifying skin conditions, the existence of racial bias against darker skin tones has been reported [1]. This racial bias may also exacerbate accuracy disparities between light and dark skin tones among non-specialists when machine learning is deployed in the clinical practice to assist physicians [5]. Addressing this issue is essential to develop and deploy safe tools in clinical practice.
Figure 1: Example predictions from our VAE model trained on a balanced subset of the Fitzpatrick17k dataset, comprising 50/50 light and dark skin tones. Notably, predictions for lighter tones are more accurate and better preserve the lesion compared to those for darker tones.
Some efforts to overcome racial bias have focused on creating large-scale dermatology image datasets that include metadata on skin tone, typically measured using the Fitzpatrick Skin Type (FST) scale. This scale goes between 1 to 6 (from the lightest to the darkest). The reference dataset to assess racial bias in dermatological images is the Fitzpatrick17k dataset, which encompasses a wide range of Fitzpatrick Skin Types (FST 1–6) [6]. While the dataset includes both skin condition and skin type labels, it remains imbalanced, with darker skin tones (FST 4–6) being underrepresented (see Fig. 2). The same study identified that existing dermatology AI models exhibit significant biases, particularly underperforming on darker skin tones. Another recent example is the PASSION dataset [4], which focuses on individuals from Sub-Saharan countries and includes FST 3–6, the most common skin type in this region. This dataset aims to address the limitations of previous datasets, which have predominantly focused on lighter skin tones.
Figure 2: Distribution of samples according to the FST in the Fitzpatrick17k dataset [6]. The label ‘-1’ represents missing values.
A different attempt to mitigate the racial bias was to generate new synthetic samples using large generative models based on diffusion (DALL·E 2) [12]. Generative models are a type of machine learning model that aims to learn the underlying structure of data, allowing them to produce new data points from this distribution. They achieved increased accuracy for underrepresented groups by using generative models to balance datasets. However, it remains unclear how racial bias may propagate through generative models. For instance, generative models are able to introduce racial bias even in a balanced scenario, see Fig. 1.
The focus of previous works has been on debiasing discriminative models for the classification of skin conditions. This fact raises the following question: How does the performance of a deep generative model, trained to generate clinical images of dermatology, vary across different skin tones? The way generative models encode and represent data is crucial for understanding their behavior in the context of racial bias. While these models can generate new synthetic examples, further study is needed to determine whether they reproduce existing biases in this context. In this study, we explore for the first time the racial bias of generative models in clinical dermatology images. We utilize a Variational Autoencoder (VAE) [9, 11], a popular generative model with an autoencoder architecture and a probability distribution in the latent space. We use the VAE to generate and reconstruct skin images across different skin types. For this purpose, we employ the Fitzpatrick17k dataset since it offers a varied FST scale, enabling us to investigate how racial bias may affect the performance of VAEs. Finally, we propose new directions toward fair generative models.
2 Methods
2.1 Deep Learning model
VAEs [9, 11] are a popular class of deep generative models. They aim to learn the distribution of a given set of images ${\mathbf{X}}=\{{\mathbf{x}}\}$ and generate new samples from this distribution $\{{\mathbf{x}}_{*}\}$. It leverages an autoencoder-like architecture combining an stochastic encoder ${\mathrm{q}}_{\phi}({\mathbf{x}}|{\mathbf{z}})$ (parametrized by $\phi$), and a stochastic decoder ${\mathrm{p}}_{\theta}({\mathbf{x}}|{\mathbf{z}})$ (parametrized by $\theta$). The word stochastic here means that the encoder and decoder are probability distributions (in contrast to deterministic autoencoders).
The encoder aims to estimate the optimal parameters of the distribution of the latent representation ${\mathbf{z}}$ given an image ${\mathbf{x}}$. Here, we use a Gaussian distribution for this, ${\mathrm{q}}_{\phi}({\mathbf{z}}|{\mathbf{x}})=\mathcal{N}(\mu_{\phi}({\mathbf{x}}),\text{diag}(\sigma_{\phi}^{2}({\mathbf{x}}))$. The decoder aims to reconstruct this image from the stochastic representation ${\mathbf{z}}$. Here, we utilize the following likelihood model for the decoder, ${\mathrm{p}}_{\theta}({\mathbf{x}}|{\mathbf{z}})=\mathcal{(}N)({\mathbf{x}}|f_{\theta}({\mathbf{z}}),{\mathbf{I}})$. Furthermore, we impose a a prior distribution to the latent variable ${\mathrm{p}}({\mathbf{z}})=\mathcal{N}({\mathbf{z}}|0,{\mathbf{I}})$. We perform variational inference to obtain the optimal variational parameters, i.e., $\{\theta,\phi\}$. For this purpose, we maximize a lower bound on the marginal log-likelihood (ELBO):
$\textrm{ELBO}(\theta,\phi;\mathbf{x})=\mathbb{E}_{q_{\phi}(\mathbf{z}|\mathbf{x})}\left[\log p_{\theta}(\mathbf{x}|\mathbf{z})\right]-\textrm{KL}(q_{\phi}(\mathbf{z}|{\mathbf{x}})||p(\mathbf{z})),$ (1)
where $\textrm{KL}(q_{\phi}(\mathbf{z}|{\mathbf{x}})||p(\mathbf{z}))$ is the Kullback-Leibler (KL) divergence between the approximate posterior $q_{\phi}(\mathbf{z}|{\mathbf{x}})$ and the prior $p(\mathbf{z})$. The first term in Eq. 1 corresponds to the negative of the reconstruction error. In this case, it is the mean square error (MSE) because of the Gaussian likelihood. The second term uses KL divergence to encourage fidelity to the prior distribution. This term can be seen as a regularizer. The training objective is to maximize the ELBO given the observational data, and the model can be optimized with respect to the variational $\phi$, and generative parameters $\theta$. To estimate the distribution ${\mathrm{p}}_{\theta}({\mathbf{z}}|{\mathbf{x}})$, we utilize Monte Carlo sampling with the reparameterization trick to obtain samples of the latent variable ${\mathbf{z}}$, i.e., ${\mathbf{z}}=\mu({\mathbf{x}})+\epsilon\odot\sigma({\mathbf{x}})$, $\epsilon\sim\mathcal{N}(0,{\mathbf{I}})$.
2.2 Perceptual loss
When working with high-fidelity images, using a standard VAE may not be the best approach. The generated images are prone to be blurry [8]. This is because using a pixel-by-pixel loss that does not capture spatial correlations between two images. A typical example is the same image translated by a few pixels. To the human eyes, they are practically the same image. However, it will result in high pixel-by-pixel loss.
This drawback of VAEs working with images has been solved by feature perceptual loss [7]. This loss between two images is defined by the difference between the hidden maps in a pretrained convolutional neural network $\Phi$. The main idea behind this approach is that the feature maps of this network have sensible information about spatial correlation in the images. Namely, it will approach better to the human vision. As a result, we can get a better visual quality output image. The loss function of the VAE with the perceptual loss used in this work is as follows,
$\mathcal{L}=\mathrm{ELBO}+\frac{1}{2C^{l}W^{l}H^{l}}\sum^{C^{l}}_{c=1}\sum^{W^{l}}_{w=1}\sum^{H^{l}}_{h=1}(\Phi({\mathbf{x}})^{l}_{c,w,h}-\Phi(f_{\theta}(x))_{c,w,h})^{2},$ (2)
where $C^{l},H^{l},W^{l}$ are the channels, height, and widht of the $l$-th feature map of the network, respectively.
2.3 Fitzpatrick17k dataset
For our study, we utilize the Fitzpatrick17k dataset [6]. To the best of our knowledge, this is the largest publicly available dataset with information on the Fitzpatrick scale. This dataset was sourced from two open-source dermatology atlases. It comprises 16,577 images with their corresponding skin condition and Fitzpatrick labels. The information on the specific skin condition is not relevant to our study since we focus on the image generation process regardless of the specific skin condition.
The Fitzpatrick skin-type labels were assigned by a team of human annotators from Scale AI. The labels were obtained by a consensus process from two to five annotators. They carried out a process to achieve the desired level of agreement, resulting in 72,277 annotations in total. The distribution of the Fitzpatrick labels of this dataset is depicted in Fig. 2. We observe that darker skin tones are notably underrepresented in the dataset, and we will need to subsample to create balanced scenarios. This information is particularly interesting for us since we focus on assessing the racial bias of the generative model.
2.4 Experimental details
Neural network architecture. The encoder and decoder of the VAE utilize residual blocks along with down-sampling and up-sampling, respectively. Each convolution block employs batch normalization and ELU activation for stable training. The resulting latent space has shape $8\times 8\times 64$. For the perceptual loss in equation (2), we used a VGG19 pretrained on ImageNet [13]. The perceptual loss was computed using the output of every convolutional layer, $16$ layers in this case.
Optimization details. We train the models for $15$ epochs after which we observed a convergence of the training loss. The optimizer was Adam with a learning rate of $10^{-4}$. The training was performed in mini-batches of $64$ samples.
Implementation. We utilize the implementation of a VAE with perceptual loss available at GitHub [^1] https://github.com/LukeDitria/CNN-VAE (accessed 2024-01-10).. The code was implemented in PyTorch 2.0.1. The code was executed using one GPU NVIDIA GeForce RTX 3090 to train and evaluate the models.
Sensitive groups. We divided the Fitzpatrick17k into two different groups. One with lighter skin, comprising FST 1-2, and one with darker skin, comprising FST 4-6. Images without a Fitzpatrick label were discarded, along with those labeled as FST 3-4, to avoid ambiguities, as these categories lie on the boundary between lighter and darker skin tones.
Experimental setting. The experiment is run ten independent times. For each run, a test set of $500$ samples is sampled from the dataset. We made two independent test sets, one with lighter images (FST 1-2) and the other with darker ones (FST 5-6). Then, with the remaining images, we sample three training set configurations (with replacement) of $1668$ samples. The ‘A’ set has 100% of lighter images, the ‘B’ set has 50/50, and the ‘C’ set has 100% of darker images.
Metrics. We compute two different metrics to assess the performance of the models in the different training and test sets. We utilize the likelihood, i.e., the MSE or reconstruction error, and the latent standard deviation given by the stochastic encoder.
3 Results
Figure 3: Likelihood or Mean Square Error (MSE) of the VAEs in the test sets.
Figure 3 shows the boxplots of the likelihood (or MSE) for our VAE with perceptual loss across the three different training set configurations. For each training set configuration, we plot the results for the two test sets: FST 1-2 and FST 5-6. The VAE generally performs well with an MSE below 0.07. This means that the neural network generally assigns a high likelihood to the data. However, the boxplot shows a performance gap between the two population subgroups. In configuration ‘A’, only trained with FST 1-2, the performance gap is quite large, implying that the reconstructed images of the VAE are biased against the darker skin tones. In this case, the lighter skin has the highest likelihood of the three possible configurations. In configuration ‘B’, when training with an equal 50/50 sample distribution from both subgroups, a significant performance gap still persists between the two groups. This finding suggests that the darker skin tones in this dataset make it more difficult to extract meaningful features for the reconstruction than the lighter skin tones. In configuration ‘C’, when training with samples of FST 5-6, the opposite occurs compared to configuration ‘A’. Here, the performance gap is not as large as in ‘A’. This corroborates the hypothesis that the generative model is generally biased against darker skin tones.
To illustrate these results, we include visualizations of the predicted reconstructions in Figs. 4 and 5. The reconstructions produced by each of the three different training sets are shown to compare the effect of the training set on performance with respect to racial bias. For lighter skin tones (see Fig. 4), the reconstructions do not show significant visual variation. However, for darker skin tones (see Fig. 5), the best reconstructions are obtained with Set C (where 100% of the samples are dark). Notably, the color tone changes as more darker examples are introduced into the training set. This observation suggests that darker skin tones in clinical skin images are more sensitive to the training set composition, and the recovery of skin tone and details of the injury is more challenging.
Figure 4: Example reconstruction of lighter skin tones. Reconstructions are produced using three training set configurations: ‘A’, ‘B’, and ‘C’.
Figure 5: Example reconstruction of darker skin tones. Reconstructions are produced using three training set configurations: ‘A’, ‘B’, and ‘C’.
Figure 6: Averaged standard deviation of the latent variable of the VAEs in the test sets.
Figure 6 depicts the values of the averaged latent standard deviation of the latent variable ${\mathbf{z}}$ for our VAE with perceptual loss in the three different configurations of the training set. This figure has been made using the same process as in the previous Fig. 3. In this case, we do not observe differences between the estimated standard deviation between the two population subgroups. Basically, the standard deviation follows a similar distribution in both cases. This suggests that the inherent mechanism of the VAE in the latent space to measure the uncertainty of the latent representation is not a proxy for the fairness of the model since the likelihood has a gap between the two subgroups.
4 Conclusion
Fairness in AI systems for healthcare is an urgent topic to ensure trustworthy AI systems are deployed in clinical practice. Clinical images of dermatology have been observed that show racial bias, as they often fail to adequately represent the wide range of skin tones present in the population. This bias has been a matter of study for the AI predictive model for computer-aided diagnosis. In this study, we show some initial results on how this bias may affect deep generative models.
In general, the VAE model proposed in this study is biased against darker skin tones, which aligns with previous findings with discriminative models [6, 5]. We have observed that the VAE assigns a lower likelihood to reconstructions of skin tones not seen during training. In general, lighter skin tones are easier to reconstruct. Even when the training set is balanced with a 50/50 distribution of the two subgroups, the VAE is still biased against darker skin tones. We have also confirmed that the latent standard deviation of the VAE does not provide any information about the fairness of the model and can not be used as a proxy for fairness.
The results obtained here suggest that the predicted uncertainty of VAEs is not adequate or informative for assessing the fairness of deep generative models. This raises concerns about potential biases when these models are deployed to generate synthetic images of skin pathologies. Future work will focus on developing enhanced uncertainty mechanisms that can effectively capture and relate to the fairness of VAEs.
{credits}
4.0.1 Acknowledgements
This work was done during Miguel López-Pérez’s visit to the Technical University of Denmark in the summer of 2024. This visit was funded by MICIU/AEI/10.13039/501100011033 and the European Union’s “NextGenerationEU”/PRTR through grant JDC2022-048318-I. The work was partly funded by the Novo Nordisk Foundation through the Center for Basic Machine Learning Research in Life Science (NNF20OC0062606). AF was supported by research grant (0087102) from the Novo Nordisk Foundation. SH was supported by a research grant (42062) from VILLUM FONDEN and from the European Research Council (ERC) under the European Union’s Horizon programme (grant agreement 101125993).
[…]
The authors have no competing interests to declare that are relevant to the content of this article.